ON THE ORIGIN OF THE BAND STRUCTURE OBSERVED IN THE COLLISION-INDUCED ABSORP-TION BANDS OF CO_{2}

Y. I. BARANOV, \underline{W. J. LAFFERTY}, AND G. T. FRASER, Optical Technology Division, NIST, Gaithersburg, MD 20899; A.A. VIGASIN, General Physics Institute, Russian Academy of Sciences, Vavilova 38, Moscow, 119991, Russia.

Two collision-induced (CIA) IR bands of CO_{2} are observed in the region of the Raman allowed ν_{1}- 2ν_{2} Fermi-dyad monomer bands in the 7 μm region. These bands consist of a featureless CIA component upon which are superimposed very distinctive CO_{2} dimer bands. The original observation of band structure in these bands was made by Welsh et al.^{a} and was interpreted by Mannik et al. ^{b} to be the P-, Q- and R-branches of a T-shaped dimer. However, molecular beam studies^{c} have subsequently shown that the dimer bands consist of nearly equally intense a-type and b-type transitions and that the dimer structure is a slipped-parallel arrangement with C_{2h} symmetry. Recently Vigasin and Baranov^{d} have modeled the dimer profile observed in room temperature CIA spectra using a symmetric-rotor model which leads to a derived C-C separation of 4.46 Å considerably larger than the molecular beam value of 3.3986 Å. In this report, we suggest an alternative explanation for the dimer band profile observed. We have modeled the bands using a Watson asymmetric-rotor Hamiltonian and the rotational constants derived in a molecular beam study. We have varied only the upper-state A rotational constant and the μ_{a}/μ_{b} ratio. In this model, the band structure is not due to P- Q, and R-branches but rather to a central sharp a-type Q-branch with broad b-type Q-branches on either side of the band center. All the Q-branch transitions sit on a pedestal of unresolved ^{p}P- and ^{r}R-branch lines. Comparison with observed spectra is good but not perfect and departures from the model will be discussed.

aH.L. Welsh, M.F. Crawford and J.L. Locke, Phys. Rev, 76, 580 (1949)

bL. Mannik, J.C. Stryland and H.L Welsh, Can. J. Phys. 49, 3056 (1971).

cK.W. Jucks et al., J. Chem. Phys., 88, 2185(1988); M.A. Walsh et al., Chem. Phys. Lett. 142, 265 (1987).

dY.I. Baranov and A.A. Vigasin, J. Mol. Spectrosc. 193, 319 (1999); A.A. Vigasin, J. Mol. Spectrosc. 200, 89 (2000).

Time required: 15 min Session in which paper is recommended for presentation: 8 Comment: This is neither jet or beam work; however it concerns the CO_{2} dimer molecule profile using jet data. We think it would be of interest to Category 8 people. If it is deemed non-suitable then please put it in category 5